A simple pendulum of length L oscillates back and forth. If this length is changed by a factor of 1.200, by what factor does the frequency of the oscillation change?
Answer #1

f1/f2 = √(L2/L1) = √1.2

f2/f1 = 1/√1.2 = 0.913 times